# Recovery speed

### v1

@@ B.2 Approximaion of $\eta^{(l)}(\infty)$
The proof of monotonically decreasing of $\rho(A^{{(l)}}_{s}(k))$ is the same as that of theorem B.4. ∎
@@ Appendix C Details of Experimential Setup
This section provides a detailed description of the experimental settings used in the experiments discussed in the main text. The experiments were conducted using the PyTorch framework.
@@ C.1.1 LSTM Setup
The model dimensions are provided in Table 1. Twenty models were generated with weights sampled uniformly from the intervals $W_{*}\in[0,0.1]$, $U_{*}\in[0,1.0]$ and $C\in[0,1.0]$
@@ C.1.2 Evaluation
We define the nominal input data $x$ and the perturbed input data $\widehat{x}$ that satisfies the assumptions of Theorem 4.3 with $t_{0}=20$ as follows.
$x(t)=0.5,\ \widehat{x}(t)=\begin{cases}1.0&\text{if }t=20,\\ 0.5&\text{else}.\end{cases}$
where $t=0,1,2,\cdots,99$. $T_{R}(x,\widehat{x})$ and $\bar{T}_{R}(k)$ were calculated according to Algorithm 1 and Algorithm 2 where $\mathcal{T}=100$. It should be noted that in cases where the model exhibits instability or $\rho(A_{s}(k))\geq 1$, this algorithm computes $T_{R}=\infty$ or $\bar{T}_{R}(k)=\infty$. Calculating statistical measures such as estimation errors or correlation coefficients using infinite values would result in arbitrary and meaningless numbers. Indeed, for some values of $k$ and specific models, our experiments showed that $\bar{T}_{R}(k)=\infty$ and $\rho(A_{s}(k))\geq 1$ occurred. This result is theoretically correct and does not indicate any deficiency in the theory. However, using infinite values to calculate statistical measures such as estimation errors or correlation coefficients would result in arbitrary and meaningless numbers. To address this issue, we replaced infinity values with finite proxies: $\bar{T}_{R}(k)=\mathcal{T}=100$ when computing these statistical measures.
@@
1:  $T\leftarrow\textrm{length}(x)$
2:  $d\leftarrow\{|y(t,s(0),x(0:t))-y(t,s(0),\widehat{x}(0:t))|\}_{t=0,1,\cdots,T}$
@@ C.2.1 LSTM Setup
Table 2: Model dimensions of two-tank system LSTM
@@ C.2.2 Data Generation
The training and test data were randomly generated through the following process Jung et al. (2023); Schoukens & Noël (2017). First, the continuous two-tank system
@@
was discretized using the Runge-Kutta method with $dt=0.01$. The system parameters are shown in Table LABEL:table:system_parameters. We randomly generated a sequence of control inputs $\{u(t)\}_{t}$, set the initial state of the system to $(h_{1}(0),h_{2}(0))=(0,0)$, and obtained the system states $\{(h_{1}(t),h_{2}(t))\}_{t}$. Here, the generation of control inputs was set to a time series length of $100,000$, with values randomly switched every $4,000$ step following a uniform distribution $\mathcal{U}(0.5,3.0)$. Using this method, we created two independent control input sequences, one for training and one for testing.
@@
$x(t)=(u(t),h_{1}(t)+w_{1}(t),h_{2}(t)+w_{2}(t))$
This addition of noise simulates real-world factors such as sensor measurement errors and environmental uncertainties. The noise distribution was set to $w_{1},w_{2}\sim\mathcal{N}(0,{0.1}^{2})$ for the training data and to $w_{1},w_{2}\sim\mathcal{N}(0,{0.01}^{2})$ for the test data. This is because while noise with a standard deviation of about $0.01$ is expected under operational conditions, intentionally introducing larger noise during the training phase aims to improve the model’s robustness and accuracy. On the other hand, the inference target data was set using the state of the next time step as follows:
$y(t)=(h_{1}(t+1),h_{2}(t+1))$
@@
The time series data are normalized to the interval $[-1,1]$ for both input $x(t)=(u(t),h_{1}(t)+w(t),h_{2}(t)+w(t))$ and output $y(t)=(h_{1}(t+1),h_{2}(t+1))$ of the LSTM before conducting inference. The equation for normalization is given below:
$x$ $=2(x-x_{min})/(x_{max}-x_{min})-1$
$y$ $=2(y-y_{min})/(y_{max}-y_{min})-1$
where $x_{min}=(0.5,0,0)$, $x_{max}=(3.0,10,10)$, $y_{min}=(0,0)$, and $y_{max}=(10,10)$.
@@ C.2.3 Train
The training dataset is split into training and validation sets, with the first $50\%$ of the sequences allocated for training and the remaining $50\%$ used for validation purposes. To reduce computational load and augment data, the training dataset is created by segmenting one length time series $50,000$ into individual sequences $49,000$, each with a length of $1,000$, using a sliding window of step size $1$. On the other hand, as data augmentation is not required for the validation dataset, we create the dataset using both window length and step size of 1,000, solely considering computational efficiency.
@@ C.2.4 Evaluation
The inference accuracy was evaluated using MAE as the metric for the test data. For the time series, we calculated the MAE, ignoring the first $10$ steps and using non-normalized values for the evaluation.
Separate from the data used for evaluating inference accuracy, we generate the nominal input data $x$ and the perturbed input data $\widehat{x}$ for the estimation of recovery time $T_{R}$ using the following steps. First, we set the time series length to $3500$ steps, the control input to $u=1.0$, and the initial water level of the system to $(h_{1}(0),h_{2}(0))=(0,0)$. Then, following the data generation process, we created $(h_{1}(t),h_{2}(t))$ and set $x(t)=(u(t),h_{1}(t),h_{2}(t))$. Then, we define $\widehat{x}$ that satisfies the assumptions of Theorem 5.3 with $t_{0}=1010$ as:
@@ D.1.1 TRAIN
For the adversarial learning, training and validation data were generated by adding pulse signals to the input data generated in Section C.2, excluding control inputs. For data with time series length $T=50,000$, we generated randomly occurring sustained pulses. Each pulse occurrence follows a Poisson process with an average rate of $\lambda=0.001$ per unit time. The total number of pulses over the entire period T is determined according to the Poisson distribution Poisson$(\lambda T)$.Each pulse has the following characteristics:
@@ D.1.2 EVALUATION
We define $\widehat{x}_{p}$ that satisfies the assumptions of Theorem 5.3 with $t_{0}=1010$ as:
$\widehat{x}_{p}(t)=\begin{cases}(u(t),h_{1}(t)+p,h_{2}(t)+p)&\text{if }1000\leq t\leq 1010,\\ x(t)&\text{else}.\end{cases}$
where $p$ was selected from the set $\{1.0,5.0,9.0\}$, and let $\widehat{y}_{p}$ represent the output of the LSTM when $\widehat{x}_{p}$ is used as the input.
@@ D.2 Results and Analysis
Baseline DD pulse $1$ DD pulse $5$ DD pulse $10$ Our

### v4

@@ B.2 Approximation of $\eta^{(l)}(\infty)$
The proof of the monotonically decreasing of $\rho(A^{{(l)}}_{s}(k))$ is the same as that of Theorem B.4. ∎
B.4 Proof of Theorem 5.3
Theorem B.6 (Upper Bound of Recovery Time).
$\sup_{s(0)\in\prod^{L}_{l=1}S^{(l)}_{k}}T_{R}(x,\widehat{x},s(0);e)\leq\min\{t_{1}\geq 0:\tilde{\beta}(2\sqrt{n_{c}}\bar{s}(k),t;k)\leq e/\|U_{y}\|\ \text{for any}\ t\geq t_{1}\},$ (163)
where $\bar{s}(k)=(\bar{s}^{(1)}(k),\dots,\bar{s}^{(L)}(k))$ and $\bar{s}^{(l)}(k)=\|(\bar{c}^{(l)}(k),\phi(\bar{c}^{(l)}(k))\bar{\sigma}^{(l)}_{o}(k))\|$.
Proof.
Denote that
$s_{1}^{(l)}=\begin{pmatrix}c_{1}^{(i)}\\ h_{1}^{(l)}\end{pmatrix}=s^{(l)}(t_{0},s(0),x(0:t_{0})),$ (164)
$s_{2}^{(l)}=\begin{pmatrix}c_{2}^{(l)}\\ h_{2}^{(l)}\end{pmatrix}=s^{(l)}(t_{0},s(0),\widehat{x}(0:t_{0}))$ (165)
for $l=1,\dots,L$. Proposition 5.2 implies that there exists $\tilde{\beta}$ such that
$\|s^{(L)}(t-t_{0},s_{1}(t_{0}),x(t_{0}:t))-s^{(L)}(t-t_{0},s_{2}(t_{0}),x(t_{0}:t))\|\leq\tilde{\beta}(\lVert s_{1}^{(1)}-s_{2}^{(1)}\rVert,\dots,\lVert s_{1}^{(L)}-s_{2}^{(L)}\rVert,t-t_{0};k)$ (166)
and $\tilde{\beta}(s^{(1)},\dots,s^{(L)},t;k)$ is increasing with respect to $s^{(1)},\dots,s^{(L)}$ for any $t$ and $k$. On the other hand, $s_{1}^{(l)},s_{2}^{(l)}\in\mathcal{S}^{(l)}(k)$ implies
$\lVert c_{1}^{(l)}-c_{2}^{(l)}\rVert$ $\leq\sqrt{n_{c}}\lVert c_{1}^{(l)}-c_{2}^{(l)}\rVert_{\infty}\leq 2\sqrt{n_{c}}\bar{c}^{(l)}(k),$ (167)
$\lVert h_{1}^{(l)}-h_{2}^{(l)}\rVert$ $\leq\sqrt{n_{c}}\lVert h_{1}^{(l)}-h_{2}^{(l)}\rVert_{\infty}\leq 2\sqrt{n_{c}}\phi(\bar{c}^{(l)}(k))\bar{\sigma}_{o}^{(l)}(k),$ (168)
and then
$\lVert s_{1}^{(l)}-s_{2}^{(l)}\rVert=\sqrt{\lVert c_{1}^{(l)}-c_{2}^{(l)}\rVert^{2}+\lVert h_{1}^{(l)}-h_{2}^{(l)}\rVert^{2}}\leq 2\sqrt{n_{c}}\bar{s}^{(l)}(k).$ (169)
Thus, we have
$\|y(t,s(0),x(0:t)-y(t,s(0),\widehat{x}(0:t)\|$ $\leq\|U_{y}\|\|s^{(L)}(t,s(0),x(0:t))-s^{(L)}(t,s(0),\widehat{x}(0:t))\|$ (170)
$=\|U_{y}\|\|s^{(L)}(t-t_{0},s_{1},x(t_{0}:t))-s^{(L)}(t-t_{0},s_{2},x(t_{0}:t))\|$ (171)
$\leq\lVert U_{y}\rVert\tilde{\beta}(2\sqrt{n_{c}}\bar{s}(k),t-t_{0};k).$ (172)
The theorem follows immediately from this inequality. ∎
@@ Appendix C Details of Experimential Setup
This section provides a detailed description of the experimental settings used in the experiments discussed in the main text. The experiments were conducted using the PyTorch framework. The random seed was set using torch.manual_seed(1234).
@@ C.1.1 LSTM Setup
The model dimensions are provided in Table 1. Twenty models were generated with weights sampled uniformly from the intervals $W_{*}\in[0,0.1]$, $U_{*}\in[0,1.0]$, and $C\in[0,1.0]$.
@@ C.1.2 Evaluation
We define the nominal input data $x$ and the perturbed input data $\widehat{x}$ that satisfy the assumptions of Theorem 4.3 with $t_{0}=20$ as follows.
$x(t)=0.5,\ \widehat{x}(t)=\begin{cases}1.0&\text{if }t=20,\\ 0.5&\text{else},\end{cases}$
where $t=0,1,2,\ldots,99$. $T_{R}(x,\widehat{x})$ and $\bar{T}_{R}(k)$ were calculated according to Algorithm 1 and Algorithm 2 where $\mathcal{T}=100$. It should be noted that in cases where the model exhibits instability or $\rho(A_{s}(k))\geq 1$, this algorithm computes $T_{R}=\infty$ or $\bar{T}_{R}(k)=\infty$. Calculating statistical measures such as estimation errors or correlation coefficients using infinite values would result in arbitrary and meaningless numbers. Indeed, for some values of $k$ and specific models, our experiments showed that $\bar{T}_{R}(k)=\infty$ and $\rho(A_{s}(k))\geq 1$ occurred. This result is theoretically correct and does not indicate any deficiency in the theory. However, using infinite values to calculate statistical measures such as estimation errors or correlation coefficients would result in arbitrary and meaningless numbers. To address this issue, we replaced infinity values with finite proxies: $\bar{T}_{R}(k)=\mathcal{T}=100$ when computing these statistical measures.
@@
[…]
@@ C.2.1 LSTM Setup
Table 2: Model dimensions of the two-tank system LSTM
@@ C.2.2 Data Generation
The training and test data were randomly generated through the following process (Jung et al.,, 2023; Schoukens and Noël,, 2017). First, the continuous two-tank system
@@
was discretized using the Runge-Kutta method with $dt=0.01$. The system parameters are shown in Table LABEL:table:system_parameters. We randomly generated a sequence of control inputs $\{u(t)\}_{t}$, set the initial state of the system to $(h_{1}(0),h_{2}(0))=(0,0)$, and obtained the system states $\{(h_{1}(t),h_{2}(t))\}_{t}$. Here, the generation of control inputs was set to a time series length of 100,000100,000100 , 000, with values randomly switched every 4,0004,0004 , 000 step following a uniform distribution $\mathcal{U}(0.5,3.0)$. Using this method, we created two independent control input sequences, one for training and one for testing.
@@
$x(t)=(u(t),h_{1}(t)+w_{1}(t),h_{2}(t)+w_{2}(t)).$
This addition of noise simulates real-world factors such as sensor measurement errors and environmental uncertainties. The noise distribution was set to $w_{1},w_{2}\sim\mathcal{N}(0,{0.1}^{2})$ for the training data and to $w_{1},w_{2}\sim\mathcal{N}(0,{0.01}^{2})$ for the test data. This is because while noise with a standard deviation of about 0.010.010.01 is expected under operational conditions, intentionally introducing larger noise during the training phase aims to improve the model’s robustness and accuracy. On the other hand, the inference target data was set using the state of the next time step as follows:
$y(t)=(h_{1}(t+1),h_{2}(t+1)).$
@@
The time series data are normalized to the interval [1,1][-1,1][ - 1 , 1 ] for both input $x(t)=(u(t),h_{1}(t)+w(t),h_{2}(t)+w(t))$ and output $y(t)=(h_{1}(t+1),h_{2}(t+1))$ of the LSTM before conducting inference. The equation for normalization is given below:
$x$ $=2(x-x_{\mathrm{min}})/(x_{\mathrm{max}}-x_{\mathrm{min}})-1,$
$y$ $=2(y-y_{\mathrm{min}})/(y_{\mathrm{max}}-y_{\mathrm{min}})-1,$
where $x_{\mathrm{min}}=(0.5,0,0)$, $x_{\mathrm{max}}=(3.0,10,10)$, $y_{\mathrm{min}}=(0,0)$, and $y_{\mathrm{max}}=(10,10)$.
@@ C.2.3 Train
[…]
@@ C.2.4 Evaluation
The inference accuracy was evaluated using MAE as the metric for the test data. For the time series, we calculated the MAE, ignoring the first 101010 steps and using non-normalized values for the evaluation.
Separate from the data used for evaluating inference accuracy, we generate the nominal input data $x$ and the perturbed input data $\widehat{x}$ for the estimation of recovery time $T_{R}$ using the following steps. First, we set the time series length to 350035003500 steps, the control input to $u=1.0$, and the initial water level of the system to $(h_{1}(0),h_{2}(0))=(0,0)$. Then, following the data generation process, we created $(h_{1}(t),h_{2}(t))$ and set $x(t)=(u(t),h_{1}(t),h_{2}(t))$. Then, we define $\widehat{x}$ that satisfies the assumptions of Theorem 5.3 with $t_{0}=1010$ as
@@ D.1.1 TRAIN
[…]
@@ D.1.2 EVALUATION
We define $\widehat{x}_{p}$ that satisfies the assumptions of Theorem 5.3 with $t_{0}=1010$ as
$\widehat{x}_{p}(t)=\begin{cases}(u(t),h_{1}(t)+p,h_{2}(t)+p)&\text{if }1000\leq t\leq 1010,\\ x(t)&\text{else},\end{cases}$
[…]
@@ D.2 Results and Analysis
Baseline DD pulse 111 DD pulse 555 DD pulse 101010 Our